What is the volume of a sample of helium that has a mass of 7.42 × 10−3 g, and a density of 0.17847 g/L?
Answer in units of mL.
1
Expert's answer
2019-03-11T08:18:24-0400

m= d*V

V = m/d = 7.42*10-3/0.17847 g/L = 0.0416L

As 1 L = 1000 mL , then V = 0.0416 L *1000mL/L = 41.6 mL

Answer: 41.6 mL